# Komaba, Meguro-ku, Tokyo 153-8902, Japan

### v2

@@ 4.1.1 Bounds on $\gamma_{\phi}$
Figure 3: The allowed region for $\Lambda_{\text{UV}}/f$ and $H/f$ as a function of ${\bar{p}_{\Lambda}}$, shown as the shaded area, is derived from the unitarity bound imposed by the purity $\gamma_{\phi}({\bar{p}_{\Lambda}})$ in de Sitter spacetime.
Then, the unitarity condition $\gamma_{\phi}\geq 0$ implies $I_{\phi}\leq 1$ with (53). For given ${\bar{p}_{\Lambda}}$, this gives a bound on the two parameters, $\Lambda_{\text{UV}}/f$ and $H/f$. Fig. 3 shows the allowed regions for ${\bar{p}_{\Lambda}}=1,0.1,0.01$. First, we find that the bounds for ${\bar{p}_{\Lambda}}=\mathcal{O}(0.1)$ imply the following two conditions:
@@
First, we notice that the singular behavior appears only when $H\neq 0$. In fact, as we discussed in Sec. 3.2, $I_{\phi}(p)$ on flat spacetime vanishes and thus the bound is trivialized in the limit $p\to 0$. See also Fig. 1. A similar singularity in the limit $p\to 0$ on de Sitter spacetime was pointed out in Pueyo:2024twm and perturbative unitarity bounds on squeezed configurations were studied. In the next subsection, we argue that such singular behaviors are peculiar to light fields in the complementary series, which are tachyonic at the superhorizon scale.
Figure 4: The allowed region for $\Lambda_{\text{UV}}/f$ and $H/f$ as a function of ${\bar{p}_{\Lambda}}$, shown as the shaded area, is derived from the unitarity bound imposed by the purity $\gamma_{\sigma}({\bar{p}_{\Lambda}})$ in de Sitter spacetime.
@@ 4.1.2 Bounds on $\gamma_{\sigma}$
As we discuss in the next subsection, this singular behavior is due to our parameter choice $m_{\sigma}=\sqrt{2}H$, for which $\sigma$ is in the complementary series. Note that the singularity survives even in the flat space limit, which is consistent with the analysis in Sec. 3.2.
@@ 4.2 More on superhorizon limit
In the previous subsection, we encountered divergence of the purity $\gamma_{\phi,\sigma}(p)$ in the superhorizon limit $p\to 0$ of the system modes. This subsection elaborates on this point to argue that this divergence is peculiar to light fields in the complementary series.
@@
×$\quad\times\left[1+e^{-\pi\mu}e^{i\alpha(\mu)}(-k\eta_{*})^{2i\mu}\right]^{-1}\left[1+e^{-\pi\mu}e^{-i\alpha(\mu)}(-k\eta_{*})^{-2i\mu}\right]^{-1}\,,$ (65)
@@ 4.3 Extension to $N$-scalar model
If we choose the modes $\phi^{\bar{I}}_{\pm{\bm{p}}}$ of the field index $\bar{I}$ and the comoving momentum ±$\pm{\bm{p}}$ as the system, the purity can be computed using the general formula (19).
@@ 5 Conclusion
In this paper, we studied perturbative unitarity bounds on the field-space curvature in de Sitter spacetime, using the momentum-space entanglement approach recently proposed in Pueyo:2024twm . We first analyzed purity in flat spacetime and showed that the UV cutoff is set by the field-space curvature, in agreement with results from the amplitude analysis. We then extended the analysis to de Sitter spacetime, where we derived unitarity bounds that involve not only the UV cutoff and the field space curvature, but also the Hubble scale. Unlike the original paper Pueyo:2024twm , our analysis was performed without taking the superhorizon limit, which allowed us to interpolate the flat space analysis and the superhorizon analysis. In particular, we derived an upper bound on the field-space curvature of the Hubble scale order, which reflects the thermal nature of de Sitter spacetime, in addition to a bound analogous to the flat space result. We also provided a detailed discussion on the superhorizon behavior of purity, interpreted in terms of the tachyonic/non-tachyonic superhorizon behavior of light/heavy fields in the complementary/principal series.
To conclude, we outline several interesting directions for future work. First, it would be worthwhile to extend our analysis to more realistic inflationary models such as Higgs inflation Bezrukov:2007ep and quasi-single field inflation Chen:2009zp , for which perturbative unitarity bounds were previously studied under the flat space approximation Lerner:2010mq ; Giudice:2010ka ; Atkins:2010yg ; Calmet:2013hia ; Barbon:2015fla ; Ema:2020zvg ; Kim:2021pbr . Second, it is important to go beyond the perturbative analysis of unitarity bounds. For this, it would be useful to investigate purity in models with phase transitions, where the UV completion is achieved in a non-perturbative manner. Finally, it would be crucial to study analyticity of purity and to perform partial wave type expansions, with the aim of formulating a bootstrap program based on entanglement measures. We hope to revisit these issues in the near future.
@@ Appendix A Bulk-to-boundary propagator in de Sitter spacetime
In the superhorizon limit $-k\eta\ll 1$, the bulk-to-boundary propagator enjoys a power-law,
@@ Appendix B Details of purity computation
${\bm{p}}$${\bm{k}}_{1}$${\bm{k}}_{2}$${\bm{k}}_{3}$$\vec{l}$$\theta_{1}$$\theta_{2}$ $\phi_{1}$${\bm{p}}$${\bm{k}}_{1}$${\bm{k}}_{2}$${\bm{k}}_{3}$$\vec{l}$$\phi_{2}$$\theta_{1}$$\theta_{2}$$x$$y$$z$

### v3

@@ 4.1.1 Bounds on $\gamma_{\phi}$
Figure 3: The allowed region for $\Lambda_{\text{UV}}/f$ and $H/f$ as a function of $\bar{p}_{\Lambda}$}, shown as the shaded area, is derived from the unitarity bound imposed by the purity $\gamma_{\phi}({\bar{p}_{\Lambda}})$ in de Sitter spacetime. The gray region indicates $\Lambda_{\text{UV}}\leq H$, which is beyond our interests.
Then, the unitarity condition $\gamma_{\phi}\geq 0$ implies $I_{\phi}\leq 1$ with (53). For given $\bar{p}_{\Lambda}$}, this gives a bound on the two parameters, $\Lambda_{\text{UV}}/f$ and $H/f$. Fig. 3 shows the allowed regions for ${\bar{p}_{\Lambda}}=1,0.1,0.01$. For a representative choice ${\bar{p}_{\Lambda}}=\mathcal{O}(0.1)$, the bounds imply the following two conditions:
@@
Figure 4: The allowed region for $\Lambda_{\text{UV}}/f$ and $H/f$: those for $p_{\mathrm{phys}}=H$ and $p_{\mathrm{phys}}=\Lambda_{\text{UV}}$ are illustrated. Their overlap satisfies the bounds for all $p_{\mathrm{phys}}$ within the range $H\leq p_{\mathrm{phys}}\leq\Lambda_{\text{UV}}$.
First, we notice that the singular behavior appears only when $H\neq 0$. In fact, as we discussed in Sec. 3.2, $I_{\phi}(p)$ on flat spacetime vanishes and thus the bound is trivialized in the limit $p\to 0$. See also Fig. 1. A similar singularity in the limit $p\to 0$ on de Sitter spacetime was pointed out in Pueyo et al. (2024) and perturbative unitarity bounds on squeezed configurations were studied. In the next subsection, we argue that such singular behaviors are peculiar to light fields in the complementary series, which are tachyonic at the superhorizon scale.
It should be emphasized that the singularity in the IR limit $p\to 0$ mentioned above does not necessarily imply breakdown of the EFT, but rather indicates breakdown of the perturbative computation of purity. Indeed, late-time correlators in this setup are IR finite, and perturbative computation works well without suffering from tachyonic behavior at the superhorizon scale. This suggests that for light fields, purity does not offer an appropriate ground to study perturbative unitarity bounds for cosmological perturbations.  [^6] Notice, however, that the unitarity bound on purity has to be respected at the non-perturbative level, even if its perturbative computation breaks down. It would be important to explore other observables more appropriate for this purpose, leaving it for future work. Nevertheless, with this caveat in mind, we expect that purity can still be used to estimate the UV cutoff scale of the EFT by choosing the system mode inside the horizon $p_{\mathrm{phys}}\gtrsim H$, for which purity is free from the IR divergence peculiar to purity of light fields. In Fig. 4, we illustrate the bounds evaluated at $p_{\mathrm{phys}}=H$ and $p_{\mathrm{phys}}=\Lambda_{\text{UV}}$ in order to identify the region of parameter space that remains allowed for all momenta within the range $H\leq p_{\mathrm{phys}}\leq\Lambda_{\text{UV}}$. There, we find that $f$ has to be bigger than $\Lambda_{\text{UV}}$ and $H$, similarly to the massive case. The same remark applies to $\gamma_{\sigma}$ as well.
@@ 4.1.2 Bounds on $\gamma_{\sigma}$
Figure 5: The allowed region for $\Lambda_{\text{UV}}/f$ and $H/f$ as a function of $\bar{p}_{\Lambda}$}, shown as the shaded area, is derived from the unitarity bound imposed by the purity $\gamma_{\sigma}({\bar{p}_{\Lambda}})$ in de Sitter spacetime. The gray region indicates $\Lambda_{\text{UV}}\leq H$, which is beyond our interests.
@@
As we discuss in the next subsection, this singular behavior is due to our parameter choice $m_{\sigma}=\sqrt{2}H$, for which $\sigma$ is in the complementary series. Note that the singularity of $I_{\sigma}$(p) persists even in the flat-space limit, which is consistent with the analysis in Sec. 3.2.
@@ 4.2 More on superhorizon limit
In the previous subsection, we encountered divergence of the purity $\gamma_{\phi,\sigma}(p)$ in the superhorizon limit $p\to 0$ of the system modes. In this subsection, we elaborate on this point and argue that this divergence is peculiar to light fields in the complementary series.
@@
$\quad\times\left[1+e^{-\pi\mu}e^{i\alpha(\mu)}(-k\eta_{*})^{2i\mu}\right]^{-1}\left[1+e^{-\pi\mu}e^{-i\alpha(\mu)}(-k\eta_{*})^{-2i\mu}\right]^{-1}\,$, (65)
@@ 4.3 Extension to $N$-scalar model
If we choose the modes $\phi^{\bar{I}}_{\pm{\bm{p}}}$ of the field index $\bar{I}$ and the comoving momentum $\pm{\bm{p}}$ as the system, the purity can be computed using the general formula (19).
@@ 5 Conclusion
In this paper, we studied perturbative unitarity bounds on the field-space curvature in de Sitter spacetime, using the momentum-space entanglement approach recently proposed in Pueyo et al. (2024). We first analyzed purity in flat spacetime and showed that the UV cutoff is set by the field-space curvature, in agreement with results from the amplitude analysis. We then extended the analysis to de Sitter spacetime, where we derived unitarity bounds that involve not only the UV cutoff and the field-space curvature, but also the Hubble scale. Unlike the original paper Pueyo et al. (2024), our analysis was performed without taking the superhorizon limit, which allowed us to interpolate the flat space analysis and the superhorizon analysis. In particular, we derived an upper bound on the field-space curvature of the Hubble scale order, which reflects the thermal nature of de Sitter spacetime, in addition to a bound analogous to the flat space result. We also provided a detailed discussion on the superhorizon behavior of purity, interpreted in terms of the tachyonic/non-tachyonic superhorizon behavior of light/heavy fields in the complementary/principal series.
To conclude, we outline several interesting directions for future work. First, it would be worthwhile to extend our analysis to more realistic inflationary models such as Higgs inflation Bezrukov and Shaposhnikov (2008) and quasi-single field inflation Chen and Wang (2010), for which perturbative unitarity bounds were previously studied under the flat space approximation Lerner and McDonald (2010); Giudice and Lee (2011); Atkins and Calmet (2011); Calmet and Casadio (2014); Barbon et al. (2015); Ema et al. (2020); Kim et al. (2021). Second, it is important to go beyond the perturbative analysis of unitarity bounds. For this, it would be useful to investigate purity in models with phase transitions, where the UV completion is achieved in a non-perturbative manner. Finally, it would be crucial to study analyticity of purity and to perform partial wave type expansions, with the aim of formulating a bootstrap program based on entanglement measures. We hope to revisit these issues in the near future.
@@ Appendix A Bulk-to-boundary propagator in de Sitter spacetime
In the superhorizon limit $k\eta\ll 1$, the bulk-to-boundary propagator enjoys a power-law,
@@ Appendix B Details of purity computation
$\bm{p}$${\bm{k}}_{1}$${\bm{k}}_{2}$${\bm{k}}_{3}$$\vec{l}$$\theta_{1}$$\theta_{2}$ $\phi_{1}$$\bm{p}$${\bm{k}}_{1}$${\bm{k}}_{2}$${\bm{k}}_{3}$$\vec{l}$$\phi_{2}$$\theta_{1}$$\theta_{2}$xxyyzz